The co efficient of thermal expansion of water is 2×10 to the power - 4 per °C. 200 gm of water at 25 °C is heated to 50 °C under 2 atm. External pressure.given density of water at 25° C is 0.9970 gm /c.c and Cp =75.30JK-1 mole - 1. Calculate W, del H, Q
1
Expert's answer
2019-08-27T04:44:39-0400

Solution.

"W = P \times \Delta V"

E - coefficient of thermal expansion of water

"E = \frac{\Delta V}{V0 \times \Delta t}"

"\Delta V = E \times V0 \times \Delta t"

"m = \rho \times V0"

"V0 = \frac{m}{\rho}"

"V0 = 200.60 cm^3"

"\Delta V = 2 \times 10^{-4} \times 200.60 \times (50-25) = 1.003 cm^3 = 0,000001003 m^3"

"W = 202650 \times 0,000001003 = 0,203 J"

"\Delta H = Cp \times \Delta T"

"\Delta H = 75.30 \times (323-298) = 1882.50 \frac{J}{mole}"

"Q = c(H2O) \times m \times \Delta t"

"Q = 4200 \times 0.2 \times 25 = 21000 J"

Answer:

"W = 0.203 J"

"\Delta H = 1882.50 \frac{J}{mole}"

"Q = 21000J"